Mrs. McTavish wants to establish an annual $9,000 scholarship in memory of her husband. The first scholarship is to be awarded two years from now. If the funds can earn 2.75% compounded annually, what amount must Mrs. McTavish pay now to sustain the scholarship in perpetuity? (Do not round intermediate calculations and round your final answer to 2 decimal places.)
Answer #1

Annual scholarship each year 2 years from now = $9000

Interest rate = 2.75% compounded annually

calculating the Present Value of Perpetuity scholarship at 1 year from now(t=1)

PV_1=\frac{Annual Scholarship}{Interest rate}

PV_1=\frac{9000}{0.0275}

PV1 = $327,272.73

Now, Calculating its Value today:-

PV1 = PV0/(1+r)^1

= $327,272.73/(1+0.0275)^1

= $318,513.60

So, the amount Mrs. McTavish must pay now to sustain the scholarship in perpetuity is $318,513.60

If you need any clarification, you can ask in comments.     

If you like my answer, then please up-vote as it will be motivating